40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 7919     Entry time: Fri Jan 18 15:08:13 2013
Author: jamie 
Type: Update 
Category: Alignment 
Subject: alignment of temporary half PRC 

[jenne, jamie]

Jenne and I got the half PRC flashing.  We could see flashes in the PRM and PR2 face cameras.

We took out the mirror in the REFL path on the AP that diverts the beam to the REFL RF pds so that we could get more light on the REFL camera.  Added an ND filter to the REFL camera so as not to saturate.

ELOG V3.1.3-